OpenStudy (anonymous):

Write the equation 3x + 2y = -6 in slope-intercept form. Then, identify the slope.

OpenStudy (anonymous):

answers could be y = 3/2x – 3, Slope is -3 y = -3/2x + 3, Slope is -3/2 y = -3x + 2, Slope is -3 y = -3/2x – 3, Slope is -3/2